OpenStudy (anonymous):

angleABD and angleBDE are supplementary. Find the measures of both angles. angleABD = 5x°, angleBDE = (17x − 18)°

jimthompson5910 (jim_thompson5910):

`angleABD and angleBDE are supplementary` so they add to 180 degrees

jimthompson5910 (jim_thompson5910):

angleABD + angleBDE = 180 5x + (17x - 18) = 180 5x + 17x - 18 = 180 22x - 18 = 180 I'll let you finish up